OpenStudy (anonymous):

Write an equation of the line through each point. Use slope-intercept form (1,-1); parallel to y=2/5x-3

OpenStudy (ajprincess):

Compare the equation y=2/5x-3 with the general equation y=mx+c where m-slope c-intercept

OpenStudy (ajprincess):

Can u find the value of m for the equation y=2/5x-3 @ChasingMaggie

OpenStudy (anonymous):

isn't the value of m 2/5?

OpenStudy (ajprincess):

yes m=2/5. for parallel lines the slopes are same. so the slope of the required line is 2/5. \[m=\frac{y-y_1}{x-x_1}\] \((x_1,y_1)\) is given as (1,-1). substituting this in the above formula solve the equation for y.

OpenStudy (ajprincess):

Does that help?

OpenStudy (anonymous):

I understand slope intercept form, but I don't understand how (1,-1) comes into play. Do I have to out that into the equation that is already laid out?

OpenStudy (ajprincess):

That is a point that lies on the required line. to calculate the slope of a line u need two points. but here u knw only one point. so u assume the other point as (x,y). u knw the slope too. Applying these values in the slope formula and solving for y will give u the required equation

OpenStudy (anonymous):

Just kidding, now I'm completely confused.

OpenStudy (anonymous):

would it be 1=2/5(-1-3)?

OpenStudy (ajprincess):